If x = (sqrt 3 + sqrt 2 ) / (sqrt 3 – sqrt 2) and y = (sqrt 3 – sqrt2 ) / ( sqrt 3 + sqrt 2 ) , find the value of x square + y square 

Grade:9

1 Answers

Arun
5 years ago
Dear student
 
on rqtionalizing , 
x = 3 + 2 + 2 sqrt 6 = 5 + 2 sqrt 6
 
similarily
y = 3 + 2 – 2 sqrt 6 = 5 – 2 sqrt 6
 
Hope it helps
 
Regards
Arun